The product of negative three and a number is twenty-four. how do you translate this into an equation and then solve?
Answer by Hawksfan(61)

The product of negative three and a number is twenty-four. how do you translate this into an equation and then solve?
Product means the result of a multiplication.
So we need to multiply
-3 * (some number) = 24
-3 * X = 24
Divide by -3 on both sides...
X = -8